OpenStudy (anonymous):

Solve cos x tan x - sin2x = 0 for all real values of x. Solve 2 sin2x – sin x = 0 for principal values of x. Write the equation 3x + 4y – 7 = 0 in normal form. Write the standard form of the equation of a line for which the length of the normal is 3 and the normal makes an angle of 135° with the positive x–axis.

OpenStudy (anonymous):

if you can explain each one and how to do them i would be grateful! :)

OpenStudy (abb0t):

HINT: \[\tan(x) = \frac{ \sin(x) }{ \cos(x) }\] \[\sin(2x)= 2\sin(x)\cos(x)\]

OpenStudy (anonymous):

im still lost .. lol

OpenStudy (abb0t):

\[x \left( \frac{ \sin(x) }{ \cos(x) } \right) -2\sin(x)\cos(x) = \frac{ xsin(x) }{ \cos(x) }-2\sin(x)\cos(x)=0\] put everything into one single fraction and solve. For the second one: \[2(2\sin(x)\cos(x))-\sin(x)=4(\sin(x)\cos(x)-\sin(x) = 0\] factor out sin(x) and find the values when sin(x)=0 and when 4cos(x)-1 = 0

OpenStudy (anonymous):

okay, thank you !